OpenStudy (anonymous):

What is the x-intercept and y-intercept for 2x - 7y = 14

OpenStudy (anonymous):

can you rewrite the equation as y = ....

OpenStudy (anonymous):

Basically an x-int means what is the value of x when y is 0 To find it, simply put in 0 in place of y in the function and solve for x as follows 2x-7(0)=14

OpenStudy (anonymous):

2x-0=14 2x=14 x=14/2 x=7 and the x-int is 7 To find the y-int its the opposite of how you find the x 2(0)-7y=14 0-7y=14 -7y=14 y=14/-7 y=-2